OpenStudy (anonymous):

If (x + 1)(x - 3) = 5, then which of the following statements is true? x + 1 = 0 or x - 3 = 0 x + 1 = 5 or x - 3 = 5 x - 4 = 0 or x + 2 = 0

OpenStudy (kropot72):

None of the statements is true. Try testing each to confirm

OpenStudy (anonymous):

x+1=5or x-3=5~~~~ should one of them be 1

OpenStudy (callisto):

(x + 1)(x - 3) = 5 x^2 -2x -3 =5 x^2 -2x -8 = 0 (x-4)(x+2)=0 (x-4)=0 or (x+2)=0

OpenStudy (anonymous):

x - 4 = 0 or x + 2 = 0

OpenStudy (anonymous):

this is true one

OpenStudy (anonymous):

yeh its a good trip

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

test it

OpenStudy (callisto):

@myko Sorry, but have you ever thought about typing the steps?

OpenStudy (anonymous):

it's evident: x - 4 = 0 or x + 2 = 0 is same as: x=4 and x=-2 plug the values in the equation to check

OpenStudy (callisto):

I'm not having doubts on your answer. But you should show your workings when you give the answer.